OpenStudy (anonymous):

How do you solve for sinx=cosx [0,2pi] ?

OpenStudy (anonymous):

\[\sin x=\cos x\\ \frac{\sin x}{\cos x}=\frac{\cos x}{\cos x}\\ \tan x=1\] When is tangent equal to 1?

OpenStudy (anonymous):

that would be at pi/4 and 5pi/4...right?

OpenStudy (anonymous):

which is kind of like asking "where is sine equal to cosine"

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

OH!!!

OpenStudy (anonymous):

funny how restating a problem makes it easier right?

OpenStudy (debbieg):

Just remember... it isn't always "safe" to divide by something like cos(x). But you can get away with it here, since the solution set requires that sin(x)=cos(x) and that never happens anywhere that cos(x)=0. :)

OpenStudy (anonymous):

yeah LOL oh, and how do I find the period of the following trig equations? y=-3tan pi x -3sec(- 6x) +2

OpenStudy (debbieg):

Just don't try it if you have: \(\Large \cos^2(x)=cos(x)\) :)

OpenStudy (anonymous):

try it?

OpenStudy (debbieg):

dividing by cos(x)

OpenStudy (anonymous):

try what? Is that impossibl?

OpenStudy (anonymous):

*impossible to solve?

OpenStudy (anonymous):

why is it bad to do that?

OpenStudy (anonymous):

the period of \[\tan(bx)\] is \(\frac{\pi}{b}\)

OpenStudy (debbieg):

If B is the coefficient on x (e.g., if you have y=sin(2x) then B=2) then: For sine, cosine, cosecant and secant, period = 2pi/B For tangent and cotangent, period = pi/B

OpenStudy (debbieg):

Sorry, @lucy4104 , did you read my first comment above also? I didn't mean to confuse you, was just saying that you can't ALWAYS go dividing an equation by an unknown (like cos(x)) but it was OK here. In a different equation, you can't do it!

OpenStudy (anonymous):

why not...?

OpenStudy (debbieg):

Because what if cos(x)=0? :)

OpenStudy (debbieg):

\[\Large \cos^2(x)=\cos(x)\]\[\Large \cos^2(x)-\cos(x)=0\]\[\Large \cos(x)(\cos(x)-1)=0\] So cos(x)=0, hence x=pi/2 or 3pi/2.... OR cos(x)=1, hence x=0 That's 3 solutions, but if you start THIS one by dividing by cos(x), you end up losing the first 2 solutions, because you eliminate the cos(x)=0 part of the solution.

OpenStudy (anonymous):

cos(x)=0... then x is pi/2 +n(pi)

OpenStudy (debbieg):

But it was OK to do so for YOUR equation above, since when cos(x)=sin(x), cos(x) is never = 0.

OpenStudy (debbieg):

Yes, if you want the general form solution set with ALL real numbers.

OpenStudy (anonymous):

oh, ok then you're saying to watch out? Wait, then how can you tell if its ok?

OpenStudy (debbieg):

In general, don't divide by an unknown if it COULD be =0. But in your original equation cos(x)=sin(x) the solution set can't include any values where cos(x)=0, since that is true (x=pi/2 and x=3pi/2 and all coterminals) ONLY at angles where cos(x) is NOT = sin(x). So you can divide by cos(x) because it can't =0. Don't get too hung up on it. :) it was kind of an aside. :)

OpenStudy (debbieg):

It's just like in algebra, if you solve \(x^2=x\) you do it by setting = 0 and factoring, not by dividing by x. Same idea.

OpenStudy (anonymous):

oh, ok now i c it~ That was a good way to put it. Thank you!